Ventura County is planning to construct a bridge across the Red Gulch to facilitate afternoon skiing in the Alpines. The first cost for the bridge will amount to $6,500,000. Annual maintenance and repairs will amount to $25,000 for each of the first five years, to $30,000 for each of the next ten years and to $35,000 for each of the next five years. In addition, a major overhaul costing $500,000 will be required at the end of the tenth year. Use an interest rate of 5% and determine the equivalent uniform annual cost for a 20-year period.
Answer #1

PW of the cost = 6500000 + 25000*(P/A, 5%, 5) + 30000*(P/A, 5%, 10)*(P/F, 5%, 5) + 35000*(P/A, 5%, 5)*(P/F, 5%, 15)

PW of the cost = 6500000 + 25000*4.329 + 30000*7.722*.7835  + 35000*4.329 *.4810

PW of the cost = $6862609

Hence,

Equivalent uniform annual cost = 6862609*(A/P, 5%, 20)

Equivalent uniform annual cost = 6862609*.0802

Equivalent uniform annual cost = $550381.24 or $550381